OpenStudy (anonymous):

How can I find a limit of ln(sin(1/x))+ln(x)?

OpenStudy (accessdenied):

What is the variable x going towards in the limit?

OpenStudy (accessdenied):

\( \displaystyle \lim_{x \to ???} \ln \left( \sin \left(1/x\right) \right) + \ln x \)

OpenStudy (anonymous):

when x approaches to infinity

OpenStudy (accessdenied):

Hm.. interesting limit. I think we could start by reducing this into terms of a single logarithm using the multiplication property: \( \log A + \log B = \log (AB) \)

OpenStudy (anonymous):

then what should i do next? i've already reached that step and got stuck

OpenStudy (accessdenied):

So you have: \( \displaystyle \lim_{x \to \infty} \ln \left( x \sin (1/x) \right) \) This looks suspiciously like the sin x / x limit as x approaches 0. Have you seen a rule about substitution like this? \( x = \dfrac{1}{u} \), \( x \to \infty \) as \(u \to 0+ \)

OpenStudy (accessdenied):

So that we have \( \displaystyle \lim_{u \to 0+} \ln \left( \dfrac{\sin u}{u} \right) \) The interior of the limit would now be approaching 1 (sin x / x limit), so the logarithm itself is approaching ln(1) = 0.